Name Date

Calculus Independent Study Path Practice Unit 9 Test

Find the following indefinite integrals:

1.

∫ tan^{−1} x dx

2.

∫ (ln x)^{2}dx

3.

∫ tan^{2} (\frac{π}{4} x) dx

4.

∫ \frac{3x − 5}{\sqrt{1 − x^{2}}} dx

5.

∫ \frac{x^{5}}{(x^{2} + 4)^{2}} dx

6.

∫ \frac{x − 1}{x^{2} + x + 1}